# A Pioneering Study and An Innovative Information Theory-based Approach to Enhance The Transparency in Phishing Detection

### v1

@@ 4.1.1. Studied datasets
We conducted the experiments on seven phishing-relevant email datasets including IWSPA-AP (i.e., the dataset was collected as part of a shared task to try and address phishing scam emails), Nazario Phishing Corpus (i.e., the dataset of phishing emails (received by one user) that are surrounded by HTML code and need to be stripped to get the useful text data), Miller Smiles Phishing Email (i.e., the dataset contains the bodies of the phishing email scams), Phish Bowl Cornell University (i.e., the dataset contains phishing emails that have been spotted and reported by students and staff at Cornell University), Fraud emails (i.e., the dataset contains fraudulent emails attempting Nigerian Letter where all the emails are in one text file and contain a large amount of header data), Cambridge (i.e., the data set contains a large number of email headers (involving information such as sending and receiving addresses and email subjects) and the body content of phishing emails), and Enron Emails (i.e., the email dataset was released as part of an investigation into Enron, which consists of emails from mostly senior management of Enron).
These datasets contain the common email header and email body. We found that most of this email data is stored in the form of web pages on public websites. We re-processed the datasets before using them for the training and testing processes in terms of removing non-ASCII characters and unrelated information to the contents of the emails such as (e.g., HTML language). We also removed duplicated or similar emails. In the end, we kept 20,000 phishing emails and 20,000 benign emails to conduct our comprehensive and rigorous experiments.
@@ 4.1.2. Data processing and embedding
We preprocessed the datasets before injecting them into deep neural networks of our AI2TALE  method and baselines. In the context of our paper, we view each email as a sequence of sentences and we aim to learn and figure out the most important sentence in each email contributing to the vulnerability label (phishing or benign) of the email.
@@
In our paper, the length of each email is padded or truncated with $L=100$ sentences (i.e., we base on the quantile values of the emails’ length of the used datasets to decide the length of each email). We observe that almost all important information relevant to the vulnerability lies in the $100$ first sentences or even lies in some very first sentences.
@@ 4.1.3. Measures
To the best of our knowledge, our method is one of the very first approaches for the phishing attack localization problem aiming to tackle and improve the explainability (transparency) of phishing detection. Therefore, studying appropriate measures for phishing attack localization is necessary and also is one of the contributions of our work. To evaluate the performance of our AI2TALE  method and the baselines in phishing attack localization, based on the explainable machine learning and email phishing domain knowledge, we introduce two main metrics to measure the quality of the most important selected sentence in each phishing-relevant email including Label-Accuracy and Cognitive-True-Positive.
@@
Via the Label-Accuracy metric, we measure whether the selected sentences obtained from each model can accurately predict the true vulnerability label (i.e., phishing or benign) of the associated email. The intuition is that the most important sentences contribute the most to the email’s vulnerability, especially to phishing emails. In our experiments, we assess each model’s top-1 selected sentence and measure if this sentence can effectively predict the email’s vulnerability without considering all sentences in the email. The higher the value of the Label-Accuracy measure, the better the model’s performance in figuring out and selecting crucial label-relevant information from the data.
@@
For the Cognitive-True-Positive metric, we investigate if the top-$1$ selected sentence successfully reflects the human cognitive principles used in phishing emails by capturing the cognitive triggers (i.e., Reciprocity, Consistency, Social Proof, Authority, Liking, and Scarcity (Akbar, 2014; Butavicius et al., 2015; Ferreira et al., 2015; Heijden and Allodi, 2019) based on the associated keywords) used in the phishing email. In particular, in our experiments, we consider the top-$1$ selected sentences of phishing emails and calculate how many percent of these selected sentences consist of cognitive triggers.
@@
It is worth noting that to measure if the top-$1$ selected sentence obtained from each method can also reflect and capture the human cognitive principles (triggers) used in each phishing email, we are particularly based on the keywords and phrases often used in each cognitive principle. To obtain the keywords and phrases, for each cognitive principle, we first base on its definition (descriptions) pioneered in the well-known work (Cialdini, 1984) (widely cited and used in phishing-related studies (Akbar, 2014; Butavicius et al., 2015; Ferreira et al., 2015; Heijden and Allodi, 2019)). We then use ChatGPT (OpenAI, 2020) to obtain all the possible keywords and phrases related to each cognitive principle’s definition (descriptions). For example, to the Scarcity principle, some of the related keywords and phrases can be relevant to Time-sensitive language (e.g., ”Act now,” or ”Expires soon”), Limited availability (e.g., ”Only a few left”), and Threats of consequences (e.g., ”Will be deleted” or ”Will lose access”). By relying on the definition (descriptions) of each cognitive principle with the help of ChatGPT in finding all the possible and main keywords and phrases, related to the definition (descriptions) of each cognitive principle, used in the Cognitive-True-Positive measure, we ensure this process is reliable and objective.
@@
Among cognitive triggers used in phishing attacks, some of them (Authority, Scarcity, and Consistency) are more popular than others (Reciprocity, Social, Proof, and Liking). To investigate how our AI2TALE  method compared to the baselines in focusing and detecting more important and popular cognitive triggers. We introduce the SAC measure, standing for the measure related to Scarcity, Authority, and Consistency of cognitive principles. In particular, we consider the top-1 selected sentences of phishing emails and calculate how many percent of these selected sentences successfully consist of the common cognitive triggers.
@@ 4.1.4. Baseline methods
The main baselines of our method are state-of-the-art interpretable machine learning approaches for phishing attack localization (i.e., where we not only aim to detect the vulnerability label (i.e., phishing or benign) of phishing-related data but also figure out and highlight the phishing information existing in the data to provide comprehensive explanations about the corresponding phishing vulnerability). It is worth noting that our proposed AI2TALE  method and the used baselines can work on the weakly supervised setting where during the training process, we only use the vulnerability label of the data and do not need to use the ground truth of phishing information in the data for the task of phishing attack localization.
Interpretable machine learning methods are strongly suitable for phishing attack localization in figuring out the most important features of the data obtained from the model to explain the model’s predictive decision. We compared the performance of our AI2TALE  method with several popular and advanced interpretable machine learning approaches including L2X ((Chen et al., 2018b)), INVASE ((Yoon et al., 2019a), ICVH ((Nguyen et al., 2021a), VIBI ((Bang et al., 2021b)), and AIM ((Vo et al., 2023)).
@@
  • L2X ((Chen et al., 2018b)). An efficient instance-wise feature selection method leverages mutual information for model interpretation. L2X aims to extract a subset of features that are most informative for each given example and the associated model prediction response.
  • INVASE ((Yoon et al., 2019a). Another effective approach for instance-wise feature selection method. INVASE consists of three neural networks including a selector network, a predictor network, and a baseline network which are used to train the selector network using the actor-critic methodology.
  • ICVH ((Nguyen et al., 2021a)). One of the first interpretable deep learning-based methods applied for source code vulnerability localization. ICVH is based on mutual information and a multi-Bernoulli distribution selection process for selecting vulnerability-relevant source code statements.
  • VIBI ((Bang et al., 2021b)). VIBI is a system-agnostic method providing a brief and comprehensive explanation by adopting an information-theoretic principle, the information bottleneck principle, as a criterion for finding such explanations.
  • AIM ((Vo et al., 2023)). A recent innovative additive instance-wise framework for model interpretation. AIM integrates both feature attribution (producing relative importance scores to each feature) and feature selection (directly identifying the subset of features most relevant to the model behavior being explained) into an effective framework for multi-class model interpretation.
@@ 4.1.5. Model’s configuration
For the L2X (Chen et al., 2018b), ICVH (Nguyen et al., 2021a), VIBI ((Bang et al., 2021b)), and AIM ((Vo et al., 2023)) methods, they were proposed to explain the output of black-box learning models. To use these methods for phishing attack localization, we keep their principles and train the models to directly approximate $p(Y\mid X)$ using $p(Y\mid\tilde{X})$ where $\tilde{X}$ consists of the selected sentences while $Y$ is ground truth label of the data $X$ instead of the output from the black-box model.
@@
We implemented our AI2TALE  method in Python using Tensorflow (Abadi et al., 2016). The trade-off parameter $\lambda$ is in $\{10^{-1},10^{-2},10^{-3}\}$ while $\sigma$ is in $\{10^{-1},2\times 10^{-1},3\times 10^{-1}\}$. For the networks $\omega\left(\cdot;\alpha\right)$ and $f\left(\cdot;\beta\right)$, we used deep feed-forward neural networks having three and two hidden layers with the size of each hidden layer in $\left\{100,300\right\}$, respectively. The dense hidden layers are followed by a ReLU function as nonlinearity and Dropout (Srivastava et al., 2014) with a retained fixed probability $p=0.8$ as regularization. The last dense layer of the network $\omega\left(\cdot;\alpha\right)$ for learning a discrete distribution is followed by a sigmoid function while the last dense layer of the network $f\left(\cdot;\beta\right)$ is followed by a softmax function for predicting. The temperature $\tau$ for the Gumbel softmax distribution is in $\{0.5,1.0\}$.
@@ 4.2. Experimental results
RQ1: Can our AI2TALE  method successfully solve the phishing attack localization in terms of predicting and giving explanations in a weakly supervised setting? Can the selected information (i.e., sentences) also successfully reflect the human cognitive principles used in phishing emails?
@@
We compare the performance of our AI2TALE  method with the baselines including L2X ((Chen et al., 2018b)), INVASE ((Yoon et al., 2019a), ICVH ((Nguyen et al., 2021a), VIBI ((Bang et al., 2021b)), and AIM ((Vo et al., 2023)) in the task of phishing attack localization (in terms of not only predicting the vulnerability of the email data but also figuring out and highlighting the phishing-relevant information existing in the data. The selected information helps provide comprehensive explanations about the vulnerability of the phishing email data for the users) using the Label-Accuracy and Cognitive-True-Positive measures.
@@
The experimental results in Table 1 show that our AI2TALE  method obtains the best performance on both Label-Accuracy and Cognitive-True-Positive compared to the baselines. In particular, our AI2TALE  method performs better than the baselines from nearly (1% to 3%) and (1% to 4%) in Label-Accuracy and Cognitive-True-Positive, respectively. The results demonstrate the effectiveness and advancement of our method for phishing attack localization in learning and figuring out the meaningful and crucial information leading to the vulnerability of the email data, especially for phishing ones, compared to the baselines.
Table 1. The performance of our AI2TALE  method and the baselines for the Label-Accuracy and Cognitive-True-Positive measures on the testing set. We denote the Cognitive-True-Positive as Cognitive-TP for short.
Methods Label-Accuracy Cognitive-TP
INVASE (Yoon et al., 2019a) 98.30% 97.20%
ICVH (Nguyen et al., 2021a) 96.72% 98.10%
L2X (Chen et al., 2018b) 98.25% 97.20%
VIBI (Bang et al., 2021b) 96.65% 94.99%
AIM (Vo et al., 2023) 98.40% 97.10%
AI2TALE  (Ours) 99.33% 98.95%
Table 2. The ground truth and the predicted vulnerability label based on the selected sentence (highlighted in yellow) of each email from our AI2TALE  method.
@@
phishing
phishing
The email content:Someone can access your paypal account, so please confirm your identity to protect your account. Start shopping faster by adding a payment method email customer paypal secure? Now check the account information that belongs to you! Why is my account access limited? Your account access has been limited for the following reason (s) we need to confirm some of your account information. Your case id, for this reason, is pp 009 536 987 252. We face a problem in the ratification of the real owner of the account and for this reason. We require your prompt attention and cooperation.”
@@
The email content: ”Bulk attention! Your discover account will close soon! Dear member, we have faced some problems with your account, so please update the account. If you do not update will be closed. To update your account, just confirm your information. (it only takes a minute). It’s easy 1. Click the link below to open a secure browser window. 2. Confirm that you’re the owner of the account, and then follow the instructions.”
@@
Note: The message from the selected sentence obtained from our method exhibits cognitive triggers commonly associated with phishing attempts used in the phishing email. In particular, the message implies a sense of urgency (concern) via ”problems with your account while ”dear member” aims to establish a connection with the recipient and imply that the message comes from a trusted source. The phrase ”please update the account” creates a sense of familiarity and consistency.
phishing
phishing
@@
The email content: ”Your account limitation. We regret to inform you that access to your account has been temporarily limited. This has been done due to several failed log in attempts. Case id ax 309 05 66 to restore your account please log in correctly. If you fail to log in correctly your account will be suspended for fraud prevention. You will be able to register again for paypal only after you authenticate your profile. We apologize for the inconvenience. This measure was taken for your protection. Paypal security team.”
@@
Note: The selected sentence conveys a sense of urgency by stating that ”access to your account has been temporarily limited.” This urgency can trigger an emotional response and prompt the recipient to take quick action without thorough consideration. Furthermore, the message’s phrasing, ”we regret to inform you,” and ”access to your account has been temporarily limited,” uses action-oriented language that urges the recipient to take immediate action to address the issue. In short, the model has successfully figured out the important sentence consisting of the cognitive principles used in the phishing email.
In addition, the experimental results on the Cognitive-True-Positive measure of our AI2TALE  method and the baselines shown in Table 1 also indicate that our method and the baselines can successfully localize and figure out the cognitive triggers used in phishing emails via the selected sentences. Compared to the baselines, our proposed AI2TALE  method obtains the best performance on the Cognitive-True-Positive measure.
In summary, in the context of explaining machine learning, particularly when dealing with complex datasets such as email phishing datasets under a weakly supervised setting, the observed improvement of (1% to 3%) for Label-Accuracy and (1% to 4%) for Cognitive-True-Positive, especially within the range approaching 99% or 100%, signifies a substantial improvement.
@@
In Table 2, we show some short phishing email samples and the top-1 selected sentences of these emails from our AI2TALE  method as well as the predicted vulnerability labels of the corresponding emails based on the selected sentences. Via these qualitative results, our method demonstrates its effectiveness in learning and figuring out the important and phishing-relevant information (i.e., sentences) from phishing emails. That helps provide useful interpretations for the phishing vulnerability prediction. Via these phishing email samples, it is shown that the cognitive triggers used in these phishing emails, which are mainly related to scarcity (i.e., creates a sense of scarcity and urgency), authority (i.e., creates a sense of trust and legitimacy), and consistency (i.e., creates a sense of familiarity and consistency) principles, are successfully represented via the selected sentences. In addition, we also observe that the most important sentences causing a specific email phishing often appear in some first positions of the email to get the users’ attention.
In conclusion for RQ1: The quantitative experimental results in Table 1 on two main measures (i.e., Label-Accuracy and Cognitive-True-Positive) and the qualitative results in Table 2 show the superiority of our AI2TALE  method in achieving high performances for phishing attack localization on the used real-world email datasets over the baselines. Our method demonstrates its effectiveness in predicting the vulnerability (phishing or benign) of emails and in figuring out the phishing information in phishing emails to give a useful explanation about the vulnerability prediction for users.
Please refer to the appendix for the additional qualitative experiments (Section 4.4) of our AI2TALE  method on other phishing email samples.
RQ2: Does the top-$1$ important selected sentences from our AI2TALE  method focus more on the commonly used and important cognitive principles (i.e., Scarcity, Consistency, and Authority) in phishing emails compared to those of the baselines?
@@
We compare the performance of our AI2TALE  method with the baselines (i.e., L2X (Chen et al., 2018b), INVASE (Yoon et al., 2019a), ICVH (Nguyen et al., 2021a), VIBI (Bang et al., 2021b), and AIM (Vo et al., 2023)) on the SCA measure where we aim to investigate the capability of our method compared to the baselines in focusing and detecting more important and popular cognitive triggers.
Table 3. The performance of our AI2TALE  method and baselines for the SAC measure on the testing set.

### v2

@@ 4.1.1. Studied datasets
We conducted the experiments on seven diverse phishing-relevant email datasets including IWSPA-AP (i.e., the dataset was collected as part of a shared task to try and address phishing scam emails), Nazario Phishing Corpus (i.e., the dataset of phishing emails (received by one user) that are surrounded by HTML code and need to be stripped to get the useful text data), Miller Smiles Phishing Email (i.e., the dataset contains the bodies of the phishing email scams), Phish Bowl Cornell University (i.e., the dataset contains phishing emails that have been spotted and reported by students and staff at Cornell University), Fraud emails (i.e., the dataset contains fraudulent emails attempting Nigerian Letter where all the emails are in one text file and contain a large amount of header data), Cambridge (i.e., the data set contains a large number of email headers (involving information such as sending and receiving addresses and email subjects) and the body content of phishing emails), and Enron Emails (i.e., the email dataset was released as part of an investigation into Enron, which consists of emails from mostly senior management of Enron).
These datasets contain the common email header and email body. We found that most of the email data are stored in the form of webpages on public websites. We re-processed the datasets before using them for the training and testing processes in terms of removing non-ASCII characters and unrelated information (e.g., HTML language syntax). We also removed duplicated emails in the used datasets as well as similar emails (when they are in the same label category, i.e., phishing or benign). In the end, we utilized around 40,000 email data samples, where half of them are phishing, from the datasets to conduct our comprehensive and rigorous experiments. Note that the email data samples used in our experiments cover a wide spectrum of writing styles, tones, and structures commonly encountered in email communications (e.g., formal business correspondences, informal personal messages, transactional emails (such as order confirmations or account notifications), marketing emails promoting products or services, and automated responses to specific actions or events).
@@ 4.1.2. Data processing and embedding
We preprocessed the datasets before injecting them into deep neural networks of our AI2TALE  method and baselines. In the context of our paper, we view each email as a sequence of sentences and aim to learn and figure out the most important sentence in each email contributing to the email’s vulnerability label (phishing or benign).
@@
In our paper, the length of each email is padded or truncated with $L=100$ sentences (i.e., we base on the quantile values of the emails’ length of the used datasets to decide the length of each email). We observe that almost all important information relevant to the phishing vulnerability lies in the $100$ first sentences or even lies in some very first sentences.
@@ 4.1.3. Measures
To the best of our knowledge, our method is one of the very first approaches proposed for solving the phishing attack localization problem aiming to tackle and improve the explainability (transparency) of phishing detection. Therefore, studying appropriate measures for phishing attack localization is necessary and also is one of the contributions of our work.
To evaluate the performance of our AI2TALE  method and the baselines in phishing attack localization, based on the explainable machine learning and email phishing domain knowledge, we introduce two main metrics to measure the quality of the most important selected sentence in each phishing-relevant email including Label-Accuracy and Cognitive-True-Positive.
@@
Via the Label-Accuracy metric, we measure whether the selected sentences obtained from each model can accurately predict the true vulnerability label (i.e., phishing or benign) of the associated email. The intuition is that the most important sentences contribute the most to the email’s vulnerability, especially to phishing emails. In our experiments, we assess each model’s top-1 selected sentence and measure if this sentence can effectively predict the email’s vulnerability without considering all sentences in the email. The higher the value of the Label-Accuracy measure, the better the model’s performance in figuring out and selecting crucial and label-relevant information from the data.
@@
For the Cognitive-True-Positive metric, we aim to investigate if the top-$1$ selected sentence from each method also reflects the human cognitive principles (triggers) used in phishing emails (i.e., Reciprocity, Consistency, Social Proof, Authority, Liking, and Scarcity (Akbar, 2014; Butavicius et al., 2015; Ferreira et al., 2015; Heijden and Allodi, 2019) based on the associated keywords). In particular, in our experiments, we consider the most important (top-$1$) selected sentences of phishing emails and calculate how many percent of these selected sentences consist of cognitive triggers.
@@
It is worth noting that to measure if the top-$1$ selected sentence obtained from each method reflects and captures the human cognitive principles (triggers) used in each phishing email, we are particularly based on the keywords and phrases often used in each cognitive principle. To obtain the keywords and phrases, for each cognitive principle, we first base on its definition (descriptions) pioneered in the well-known work (Cialdini, 1984) (widely cited and used in phishing-related studies (Akbar, 2014; Butavicius et al., 2015; Ferreira et al., 2015; Heijden and Allodi, 2019)). We then use ChatGPT (OpenAI, 2020) to obtain all the possible keywords and phrases related to each cognitive principle’s definition (descriptions). For example, to the Scarcity principle, some of the related keywords and phrases can be relevant to Time-sensitive language (e.g., ”Act now,” or ”Expires soon”), Limited availability (e.g., ”Only a few left”), and Threats of consequences (e.g., ”Will be deleted” or ”Will lose access”). By relying on the definition (descriptions) of each cognitive principle with the help of ChatGPT in finding all the possible and main keywords and phrases, related to the definition (descriptions) of each cognitive principle, used in the Cognitive-True-Positive measure, we ensure this process is reliable and objective.
@@
Among cognitive triggers used in phishing attacks, some of them (Authority, Scarcity, and Consistency) are more popular than others (Reciprocity, Social, Proof, and Liking). To investigate how our AI2TALE  method compared to the baselines in consisting of more important and popular cognitive triggers via the most important selected information. We introduce the SAC measure, standing for the measure related to Scarcity, Authority, and Consistency of cognitive principles. In particular, we consider the top-1 selected sentences of phishing emails and calculate how many percent of these selected sentences consist of the commonly used cognitive triggers.
@@ 4.1.4. Baseline methods
The baselines of our method are recent, popular, and state-of-the-art interpretable machine learning approaches falling into the category of intrinsic interpretable models including L2X (Chen et al., 2018a), INVASE (Yoon et al., 2019a), ICVH (Nguyen et al., 2021a), VIBI (Bang et al., 2021b), and AIM (Vo et al., 2023) that we apply to solve the phishing attack localization problem (i.e., where we not only aim to detect the vulnerability label (phishing or benign) of phishing-related data but also learn and figure out the most important and phishing information existing in the data to provide useful and concise explanations about the corresponding phishing vulnerability). Intrinsic interpretable machine learning techniques, a.k.a. self-explanatory models, are strongly suitable for phishing attack localization because they are able to not only make predictions themselves but also figure out the most important features of the data obtained from the model to explain the model’s predictive decision (i.e., please refer to section (2) for our brief description about intrinsic interpretable techniques).
@@
  • L2X (Chen et al., 2018a). An efficient instance-wise feature selection method leverages mutual information for model interpretation. L2X aims to extract a subset of features that are most informative for each given example and the associated model prediction response.
  • INVASE (Yoon et al., 2019a). Another effective approach for instance-wise feature selection. INVASE consists of three neural networks including a selector network, a predictor network, and a baseline network which are used to train the selector network using the actor-critic methodology.
  • ICVH (Nguyen et al., 2021a). One of the first interpretable deep learning-based methods applied for source code vulnerability localization. ICVH is based on mutual information and a multi-Bernoulli distribution selection process for selecting vulnerability-relevant source code statements.
  • VIBI (Bang et al., 2021b). VIBI is a system-agnostic method providing a brief and comprehensive explanation by adopting an information-theoretic principle, the information bottleneck principle, as a criterion for finding such explanations.
  • AIM (Vo et al., 2023). A recent innovative additive instance-wise framework for model interpretation. AIM integrates both feature attribution (producing relative importance scores to each feature) and feature selection (directly identifying the subset of features most relevant to the model behavior being explained) into an effective framework for multi-class model interpretation.
It is worth noting that our proposed AI2TALE  method and the used baselines can work on the weakly supervised setting where during the training process, we only utilize the vulnerability label of the data, without requiring the ground truth of phishing information in the data for the task of phishing attack localization.
@@ 4.1.5. Model’s configuration
For the L2X (Chen et al., 2018a), ICVH (Nguyen et al., 2021a), VIBI (Bang et al., 2021b), and AIM (Vo et al., 2023) methods, they were proposed to explain the output of black-box learning models. To use these methods for phishing attack localization, we keep their principles and train the models to directly approximate $p(Y\mid X)$ using $p(Y\mid\tilde{X})$ where $\tilde{X}$ consists of the selected sentences while $Y$ is ground truth label of the data $X$ instead of the output from the black-box model.
@@
We implemented our AI2TALE  method in Python using Tensorflow (Abadi et al., 2016). The trade-off parameter $\lambda$ is in $\{10^{-1},10^{-2},10^{-3}\}$ while $\sigma$ is in $\{10^{-1},2\times 10^{-1},3\times 10^{-1}\}$. For the networks $\omega\left(\cdot;\alpha\right)$ and $f\left(\cdot;\beta\right)$, we used deep feed-forward neural networks having three and two hidden layers with the size of each hidden layer in $\left\{100,300\right\}$, respectively. The dense hidden layers are followed by a ReLU function as nonlinearity and Dropout (Srivastava et al., 2014) with a retained fixed probability $p=0.8$ as regularization. The last dense layer of the network $\omega\left(\cdot;\alpha\right)$ for learning a discrete distribution is followed by a sigmoid function while the last dense layer of the network $f\left(\cdot;\beta\right)$ is followed by a softmax function for predicting. The temperature $\tau$ for the Gumbel softmax distribution is in $\{0.5,1.0\}$. Note that we utilized the commonly used values for these hyper-parameters.
@@ 4.2. Experimental results
RQ1: Can our AI2TALE  method solve the phishing attack localization by predicting and giving explanations in a weakly supervised setting? Does the selected information (i.e., sentences) also reflect the human cognitive principles used in phishing emails?
@@
We compare the performance of our AI2TALE  method with the baselines including L2X (Chen et al., 2018a), INVASE (Yoon et al., 2019a), ICVH (Nguyen et al., 2021a), VIBI (Bang et al., 2021b), and AIM (Vo et al., 2023) in the task of phishing attack localization (in terms of not only predicting the vulnerability of the email data but also learning and figuring out the most important (top-1) and phishing-relevant information existing in the data. The selected information helps provide useful and concise explanations about the vulnerability of the phishing email data for the users) using the Label-Accuracy and Cognitive-True-Positive measures.
@@
The experimental results in Table 1 show that our AI2TALE  method obtains the best performance on both Label-Accuracy and Cognitive-True-Positive compared to the baselines. Importantly, our method achieves a significantly higher performance, with improvements ranging from approximately 1.5% to 3.5% compared to state-of-the-art baselines, measured by the combined average performance of two main metrics, i.e., Label-Accuracy and Cognitive-True-Positive. The results demonstrate the effectiveness and advancement of our method for phishing attack localization in learning and figuring out the most meaningful and crucial information leading to the vulnerability of the email data, especially for phishing ones, compared to the baselines.
Table 1. The performance of our AI2TALE  method and the baselines for the Label-Accuracy (Label-Acc) and Cognitive-True-Positive (Cognitive-TP) measures, as well as their combined average results (denoted as Average), on the testing set (the best results in bold).
Methods Label-Acc Cognitive-TP Average
INVASE (Yoon et al., 2019a) 98.30% 97.20% 97.75%
ICVH (Nguyen et al., 2021a) 96.72% 98.10% 97.41%
L2X (Chen et al., 2018a) 98.25% 97.20% 97.73%
VIBI (Bang et al., 2021b) 96.65% 94.99% 95.82%
AIM (Vo et al., 2023) 98.40% 97.10% 97.75%
AI2TALE  (Ours) 99.33% 98.95% 99.14%
$\uparrow$ ($\sim$(∼ (1.5% $\rightarrow$3.5%)
In addition, the experimental results on the Cognitive-True-Positive measure of our AI2TALE  method and the baselines shown in Table 1 also indicate that the most important (top-1) selected sentences from our method and the baselines also reflect the cognitive triggers used in phishing emails. Compared to the baselines, our proposed AI2TALE  method achieves the best performance on the Cognitive-True-Positive measure.
Table 2. The ground truth and the predicted label based on the most important (top-1) selected sentence (highlighted in yellow) of each email from our AI2TALE  method.
@@
phishing phishing The email:Someone can access your paypal account, so please confirm your identity to protect your account. Start shopping faster by adding a payment method email customer paypal secure? Now check the account information that belongs to you! Why is my account access limited? Your account access has been limited for the following reason (s) we need to confirm some of your account information. Your case id, for this reason, is pp 009 536 987 252. We face a problem in the ratification of the real owner of the account and for this reason. We require your prompt attention and cooperation.”
@@
phishing phishing The email: ”Bulk attention! Your discover account will close soon! Dear member, we have faced some problems with your account, so please update the account. If you do not update will be closed. To update your account, just confirm your information. (it only takes a minute). It’s easy 1. Click the link below to open a secure browser window. 2. Confirm that you’re the owner of the account, and then follow the instructions.”
@@
Note: The message from the selected sentence obtained from our method exhibits cognitive triggers commonly associated with phishing attempts used in the phishing email. In particular, it implies a sense of urgency (concern) via ”problems with your account while ”dear member” aims to establish a connection with the recipient and imply that the message comes from a trusted source. The phrase ”please update the account” creates a sense of familiarity and consistency.
@@
phishing phishing The email: ”Your account limitation. We regret to inform you that access to your account has been temporarily limited. This has been done due to several failed log in attempts. Case id ax 309 05 66 to restore your account please log in correctly. If you fail to log in correctly your account will be suspended for fraud prevention. You will be able to register again for paypal only after you authenticate your profile. We apologize for the inconvenience. This measure was taken for your protection. Paypal security team.”
@@
Note: The selected sentence conveys a sense of urgency by stating that ”access to your account has been temporarily limited.” This urgency can trigger an emotional response and prompt the recipient to take quick action without thorough consideration. Furthermore, the message’s phrasing, ”we regret to inform you,” and ”access to your account has been temporarily limited,” uses action-oriented language that urges the recipient to take immediate action to address the issue. In short, the model has successfully figured out the important sentence containing the cognitive principles in the phishing email.
benign phishing The email: ”Lower your mortgage payment! Bad credit, no problem! Lower your mortgage payment! Bad credit, no problem! Whether your credit is excellent or less than perfect, loanweb has a lender that can help you! Lowest rates on the web! Bad credit? Refinance to get cash! Lower monthly payments! Bad credit? Refinance to consolidate bills! Click here for lower mortgage payment! Win a free bread maker! Win free bread makers, toaster ovens, cookware and more get a free subscription to cooking pleasures magazine get a free multi purpose grater test and keep free cooking products get free recipes from world famous chefsplus get a free 90 day membership in the cooking club of america. Home equity loans without perfect credit! Click here for a home equity loan! Free application! Home equity loans up to 125 potential tax deductible interest! Customized, competitive equity lines and loans. Apply now!”
Note: In the used datasets, the ground truth label for this email is benign. However, our method predicts it as phishing due to the detection of potential phishing indicators, such as the use of enticing offers of ”lowering mortgage payments” and the presence of ”suspicious links”. These elements may prompt users to provide personal information and bank account details via the phrase ”Click here for lower mortgage payment!”. Although the email’s ground truth classification was incorrect, we deem the predicted label and the highlighted sentence from our model valuable in alerting users to potential phishing attempts.
Note that under a weakly supervised setting, when dealing with complex data (e.g., emails written from various writing styles and structures) and only the most important (top-1) selected sentences from emails are utilized, the observed improvement of our AI2TALE method from around 1.5% to 3.5% in the combined average performance of the Label-Accuracy and Cognitive-True-Positive measures, especially within the range over 99% and approaching 100%, signifies a substantial advancement.
@@
To demonstrate the effectiveness of our AI2TALE  method in solving the phishing attack localization problem, in Table 2, we present various email samples alongside the most important (top-1) selected sentences extracted by our method as well as the predicted labels of the corresponding emails based on the selected sentences. Via these qualitative results, our method showcases its effectiveness in learning and figuring out the most important and phishing-relevant information (i.e., sentences) from phishing emails. That helps provide useful and concise interpretations for the phishing prediction. Via these phishing email samples, it is evident that the commonly used cognitive triggers in the phishing emails related to scarcity (i.e., creates a sense of scarcity and urgency), authority (i.e., creates a sense of trust and legitimacy), and consistency (i.e., creates a sense of familiarity and consistency) principles, are also represented via the selected sentences. In addition, we also observe that the most important (top-1) sentence causing a specific email phishing tends to appear in some first positions of the email to get the users’ attention.
In summary, the quantitative results in Table 1 demonstrate the high quality of our AI2TALE  method for phishing attack localization, as evidenced by the Label-Accuracy and Cognitive-True-Positive measures. Additionally, the qualitative findings presented in Table 2 underscore the effectiveness of our AI2TALE  method in providing concise yet important explanations about phishing attacks. The most important (top-1) selected sentences from the emails offer valuable insights into the nature of the attacks.
In addition, in Table 2, we also present an example where our AI2TALE  method incorrectly predicts the ground truth of the email. Specifically, the ground truth of this email is labeled as benign. However, our AI2TALE  method predicts it as phishing. Through our observations and the top-1 selected sentence obtained from our method, we believe that the model has identified some potential phishing attempts within the email, i.e., the use of enticing offers of ”lowering mortgage payments” and the presence of ”suspicious links”. Therefore, despite the incorrect classification of the email’s ground truth, we consider the predicted label and the highlighted sentence from our model valuable in alerting users to possible phishing threats. Leveraging the predicted label and selected sentence, users can take proactive measures, such as contacting the loanweb directly using a verified phone number or email address, to verify the message’s authenticity.
In conclusion for RQ1: The quantitative experimental results in Table 1 on two main measures (i.e., Label-Accuracy and Cognitive-True-Positive) and the qualitative results in Table 2 show the superiority of our AI2TALE  method in achieving high performances for phishing attack localization on the real-world email datasets over the baselines. Our method demonstrates its effectiveness in predicting the vulnerability label (i.e., phishing or benign) of emails as well as in figuring out the most important and phishing-relevant information in the emails to give a useful and concise explanation about the vulnerability prediction for users.
Please refer to the appendix for the additional qualitative experiments (Section 9.2) of our AI2TALE  method on other phishing-related email samples.
RQ2: Does the most important (top-$1$) selected sentences from our AI2TALE  method primarily consist of the commonly used and important cognitive principles (i.e., Scarcity, Consistency, and Authority) in phishing emails compared to those of the baselines?
@@
We compare the performance of our AI2TALE  method with the baselines (i.e., L2X (Chen et al., 2018a), INVASE (Yoon et al., 2019a), ICVH (Nguyen et al., 2021a), VIBI (Bang et al., 2021b), and AIM (Vo et al., 2023)) on the SCA measure where we aim to investigate if our AI2TALE  method consists of more important and commonly used cognitive triggers in the most important (top-1) selected sentences from corresponding emails than the baselines.
Table 3. The performance of our AI2TALE  method and the baselines for the SAC measure on the testing set (the best results in bold).